A metal alloy weighing 2 mg and containing 45% gold is melted and mixed with 3 mg of a different alloy which contaians 30% gold. Which percent of the resulting alloy is gold.
Thanks!!!!

Answer by flame8855(424)

amount of gold in the first part :2*0.45=0.9 mg
amount of gold in the second part : 3*0.3=0.9 mg
total gold= 1.8 mg
%=1.8/5 *100 = 36 %